OpenStudy (anonymous):

if the total weight of an apple is 4/5 of it's weight plus 4/5 of an ounce,what is it's weight in ounces?

OpenStudy (anonymous):

i didnt get you plz explain properly

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

w=4/5 w+4/5 w-4/5 w=4/5 1/5 w=4/5 find w=___?

OpenStudy (anonymous):

@ laraib